# SOMA: a novel sampler for exchangeable variables

### v1

@@ Appendix A SOMA as a multi-proposal MCMC sampler
A.1 Alternative proof of Theorem 5
SOMA simultaneously considers $n$ proposed states $[x_{-1},y]$, $[x_{-2},y],\ldots,[x_{-n},y]$, with (unnormalized) weights $w_{i}(y,x).$ In this view, SOMA is indeed a multi-proposal MCMC method. The Markov transition kernel of SOMA $P_{\mathrm{SOMA}}(\cdot\mid x)$ is a special case of $P^{(K)}(\cdot\mid x)$ of Pozza and Zanella, (2025) with $K=n$. SOMA has a special way of constructing the multiple proposals and a weight function to choose among the proposed states.
In SOMA, the ‘marginal proposal’ for each component $i$ is in fact the IMH proposal $q(\cdot)$. Let
@@ A.1 Alternative proof of Theorem 5
denote the IMH proposal kernel for the $i$-th component, defined on the full state space $\mathbb{X}^{n}.$ Then $\widetilde{Q}(\cdot\mid x)=\frac{1}{n}Q_{i}(\cdot\mid x)$ is the overall proposal kernel for Ran-IMwG: it proposes an index $i$ uniformly and proposes a new from $\mathbb{X}$. We denote the Markov transition kernel associated with proposal $\widetilde{Q}$ and acceptance probability (4) by $\widetilde{P}.$ It coincides with the Markov transition kernel of Ran-IMwG, denoted $P_{\mathrm{RAN}}(\cdot\mid x).$ In other words, Ran-IMwG is the ‘single proposal’ kernel that we should compare SOMA with. Pozza and Zanella, (2025)[Theorem 1] states that
@@
which is equivalent to (10) in Theorem 5.
A.2 Asymptotic variance of SOMA and computation cost
@@
Luckily, computing the weights in parallel can reduce the runtime cost of SOMA. Oftentimes, the target distribution $\pi$ has the structure to make this parallel/vectorized implementation feasible. SOMA can thus have a computational complexity comparable to Ran-IMwG and Sys-IMwG under moderate values of $n$, using a vectorized implementation. The costs would have the same order under a parallel implementation. These allow one to harness the efficiency gained from SOMA’s design. We give Example 5, which continues with the perturbed summary statistics privacy mechanism in Examples 2 and 4.
@@
Privacy mechanisms that add data-independent noise to a query of the form $s(x)=\sum_{i=1}^{n}t(x_{i})$ have densities $\eta(s_{\mathrm{dp}}\mid x)=g(s_{\mathrm{dp}},\sum_{i=1}^{n}t_{i}(x_{i}))$ for tractable and known functions $g$ and $t$. Ju et al., (2022, Assumption 2) coined the term ‘record additivity’ to describe how $\eta(s_{\mathrm{dp}}\mid x)$ depends on an additive summary of $x$, and connects this property with implementation efficiency of data augmentation methods. With the widely used Laplace mechanism $s_{\mathrm{dp}}\sim\mathrm{Laplace}(s(x),{\Delta}/{\epsilon})$ and the weights $w_{i}=\eta(s_{\mathrm{dp}}\mid x_{-i},y)$ in Example 4, we should have
@@
The swap-one-out structure in (27) enables parallel or vectorized implementation, which can reduce SOMA’s computational complexity to a large extent. We include empirical run time comparisons in Section C.3, showing the runtime cost of SOMA is less than twice that of the Ran/Sys-IMwG methods with $n=100$ components. In this case, instead of comparing the asymptotic variance of using $T$ iterations of SOMA with $nT$ iterations of Ran-IMwG, we should use $2T$ iterations of Ran-IMwG.
Appendix B Additional proofs and technical results
B.1 Proofs of Section 3
  • Proof of Theorem 2.
    To verify detailed balance, we should derive the ‘backward proposal kernel’ $K(x\mid x^{\prime})$. In this proposal kernel, the current state is $x^{\prime}$, the proposed point is $x_{1}$, the index $i$ is selected according to weights $w_{i}(x_{1},x^{\prime})$ defined in (5) & (7). Notice that from $x^{\prime}$, we can only move to $[x^{\prime}_{-1},x_{1}]=x$ or $[x^{\prime}_{-i},x_{1}]$ which is equivalent to $[x_{-i},y]$ for $i\geq 2$ up to shuffling indices, in terms of having equal density under $\pi$. Now we have
    $w_{1}(x_{1},x^{\prime})$ $=\frac{\pi([x^{\prime}_{-1},x_{1}])}{q(x_{1})\prod_{j=2}^{n}q(x_{j}^{\prime})}=\frac{\pi(x)}{\prod_{j=1}^{n}q(x_{j})}=w_{0}(y,x),$
    $w_{k}(x_{1},x^{\prime})$ =$\frac{\pi([x^{\prime}_{-k},x_{1}])}{q(x_{1})q(y)\prod_{2\leq j\leq n,j\not=k}q(x^{\prime}_{j})}=\frac{\pi([x_{-k},y])}{q(y)\prod_{1\leq j\leq n,j\not=k}q(x_{j})}=w_{k}(y,x),\quad\text{for }k\geq 2.$
    Using these weights, the backward proposal kernel density is
    $K(x\mid x^{\prime})=q(x_{1})\frac{w_{1}(x_{1},x^{\prime})}{\sum_{k=1}^{n}w_{k}(x_{1},x^{\prime})}=q(x_{1})\frac{w_{0}(y,x)}{w_{0}(y,x)+\sum_{k=2}^{n}w_{k}(y,x)}.$
    We also have
    $\frac{\pi(x^{\prime})K(x\mid x^{\prime})}{\pi(x)K(x^{\prime}\mid x)}=\frac{\pi(x^{\prime})q(x_{1})\frac{w_{0}(y,x)}{w_{0}+\sum_{j=2}^{n}w_{j}(y,x)}}{\pi(x)q(y)\frac{w_{1}}{\sum_{j=1}^{n}w_{j}(y,x)}}=\frac{\sum_{j=1}^{n}w_{j}(y,x)}{w_{0}+\sum_{j=2}^{n}w_{j}(y,x)},$ (28)
    since
    $\pi(x^{\prime})q(x_{1})w_{0}(y,x)=\pi(x^{\prime})q(x_{1})\frac{\pi(x)}{\prod_{j=1}^{n}q(x_{j})}=\pi(x)\frac{q(y)}{q(y)}\frac{\pi([x_{-1},y])}{\prod_{j=2}^{n}q(x_{j})}=\pi(x)q(y)w_{1}(y,x).$
B.2 Proofs of Section 4.1
@@ B.1 Proofs of Section 3
    Based on Lemma 14, integrating both sides over $q(y)$ yields
    $\int_{\mathbb{X}}q(y)\sum_{i=1}^{n}\left(\frac{w_{i}(y,x)}{W(y,x)}\right)\cdot\alpha_{i}^{\mathrm{SOMA}}(y,x)\mathrm{d}y\geq\int_{\mathbb{X}}q(y)\sum_{i=1}^{n}\left(\frac{1}{n}\right)\cdot\alpha_{i}^{\mathrm{IMwG}}(y,x)\mathrm{d}y,$
@@
    Since $\varphi(u)=\mathbb{I}_{x}(u)$, the integral reduces to $[1-A(x)]\cdot\int_{\mathbb{X}^{n}}\varphi(u)\mathrm{d}P^{m-1}(u\mid x)$. By induction, we have
@@
    We first prove that $A^{\mathrm{SOMA}}(x)\geq{n}/({n+M-1})$. Let $T_{i}(y,x)=\sum_{j=1,j\neq i}^{n}w_{j}(y,x)$. Under Assumption 2, we have $w_{i}(y,x)/w_{0}(y,x)\geq 1/M$, and $T_{i}(y,x)/w_{0}(y,x)\geq(n-1)/M$ holds for all $x$ and $y$. Now the overall acceptance probability $A^{\mathrm{SOMA}}(x)$ for SOMA has the form
    $A^{\mathrm{SOMA}}(x)$ $=\int_{\mathbb{X}}q(y)\sum_{i=1}^{n}\left[\mathbb{P}(I=i)\min\left\{1,\frac{W(y,x)}{W(y,x)+w_{0}(y,x)-w_{i}(y,x)}\right\}\right]\mathrm{d}y$
    $=\int_{\mathbb{X}}q(y)\sum_{i=1}^{n}\left[\mathbb{P}(I=i)\min\left\{1,\frac{T_{i}(y,x)+w_{i}(y,x)}{T_{i}(y,x)+w_{0}(y,x)}\right\}\right]\mathrm{d}y$
    $=\int_{\mathbb{X}}q(y)\sum_{i=1}^{n}\left[\mathbb{P}(I=i)\min\left\{1,1-\frac{1-w_{i}(y,x)/w_{0}(y,x)}{1+T_{i}(y,x)/w_{0}(y,x)}\right\}\right]\mathrm{d}y$
    $\geq\int_{\mathbb{X}}q(y)\sum_{i=1}^{n}\left[\mathbb{P}(I=i)\min\left\{1,1-\frac{1-1/M}{1+(n-1)/M}\right\}\right]\mathrm{d}y$
    $=1-\frac{1-1/M}{1+(n-1)/M}$
    $=\frac{n}{n+M-1}.$ (39)
@@
    $A^{\mathrm{RAN}}(x)$ $=\int_{\mathbb{X}}q(y)\sum_{i=1}^{n}\left[\mathbb{P}(I=i)\min\left\{1,\frac{w_{i}(y,x)}{w_{0}(y,x)}\right\}\right]\mathrm{d}y$
    $\geq\int_{\mathbb{X}}q(y)\sum_{i=1}^{n}\left[\mathbb{P}(I=i)\min\left\{1,\frac{1}{M}\right\}\right]\mathrm{d}y$
    $=\frac{1}{M},$ (40)
@@
    $A_{i}^{\mathrm{SYS}}(x)$ $=\int_{\mathbb{X}}q(y)\min\left\{1,\frac{w_{i}(y,x)}{w_{0}(y,x)}\right\}\mathrm{d}y$
    $\geq\int_{\mathbb{X}}q(y)\min\left\{1,\frac{1}{M}\right\}\mathrm{d}y$
    $=\frac{1}{M}.$ (41)
@@
B.3 The coupling kernel of SOMA in Section 4.2
We present Algorithm 2 in more detail. Assume we have two chains $\Phi$ and $\tilde{\Phi}$ with corresponding states $x^{(t)}$ and $\tilde{x}^{(t)}$ at time $t$. Here is a coupling kernel.
1. Propose the same single offer $y\sim q(y)$ for both chains $\Phi$ and $\tilde{\Phi}$.
@@ 3. Compute the component-wise minimum of selection probabilities and residuals:
Figure 10: Maximal coupling of $\mathrm{Multinomial}(w)$ and $\mathrm{Multinomial}(\widetilde{w})$.
@@ 4. Calculate the acceptance probability
    Let the set of coupled indices be =$\mathcal{I}=\{i:x_{i}=\tilde{x}_{i}\}$. In the coupling construction, if $I$ and $\tilde{I}$ are both on uncoupled positions and both chains accept the proposed component $y$, then the distance between the chains will reduce by 1. This argument leads to
    $\quad P_{\mathrm{SOMA-coupled}}\left(d(x^{\prime},\tilde{x}^{\prime})=d(x,\tilde{x})-1\mid x,\tilde{x}\right)$
    $=\int_{\mathbb{X}}\sum_{i\in\mathcal{I}^{c}}\sum_{j\in\mathcal{I}^{c}}\mathbb{P}(I=i,\tilde{I}=j\mid x,\tilde{x},y)\cdot\min(\alpha_{i},\tilde{\alpha}_{j};x,\tilde{x},y)q(y)\mathrm{d}y$
    $\geq\frac{n}{n+M-1}\cdot\mathbb{P}(I\in\mathcal{I}^{c},\tilde{I}\in\mathcal{I}^{c}\mid x,\tilde{x},y)$
    $=\frac{n}{n+M-1}\cdot\min\left(\frac{\sum_{i\in\mathcal{I}^{c}}w_{i}}{\sum_{i=1}^{n}w_{i}},\frac{\sum_{i\in\mathcal{I}^{c}}\tilde{w_{i}}}{\sum_{i=1}^{n}\tilde{w}_{i}}\right)$
    $\geq\frac{n}{n+M-1}\cdot\frac{d(x,\tilde{x})}{d(x,\tilde{x})+M(n-d(x,\tilde{x}))}.$
    The first inequality uses Theorem 9. The second lower bound is achieved when ${w_{j}}/{w_{i}}=M$ for all $j\in\mathcal{I}$ and all $i\in\mathcal{I}^{c}$. We also note that the two index sets have cardinality $d(x,\tilde{x})$ and $n-d(x,\tilde{x})$ respectively.
@@
    $P_{\mathrm{Ran-coupled}}\left(d(x^{\prime},\tilde{x}^{\prime})=d(x,\tilde{x})-1\mid x,\tilde{x}\right)\geq\inf_{i\in\mathcal{I}^{c}}\left[\alpha_{i}^{\mathrm{IMwG}}(x,y)\right]\cdot\mathbb{P}(I=\tilde{I}\in\mathcal{I}^{c})=\frac{1}{M}\cdot\frac{d(x,\tilde{x})}{n}.$
@@
For any current state $(x,\tilde{x})$, let $(x^{\prime},\tilde{x}^{\prime})$ be samples from the coupled SOMA kernel $P_{\mathrm{SOMA-coupled}}$ in Algorithm 2. Under Assumption 2, we must have
$\quad P_{\mathrm{SOMA-coupled}}\left(d(x^{\prime},\tilde{x}^{\prime})=d(x,\tilde{x})+1\mid x,\tilde{x}\right)$
$\leq\frac{(n-d)M}{(n-d)M+d}-\frac{n}{n+M-1}\cdot\frac{(n-d)}{(n-d)+dM},$ (44)
where we use $d=d(x,\tilde{x})$ for simplicity.
@@
$P_{\mathrm{SOMA-coupled}}\left(d(x^{\prime},\tilde{x}^{\prime})=d(x,\tilde{x})+1\mid x,\tilde{x}\right)\leq\frac{M(M-1)}{(M+1)^{2}}.$ (45)
@@
    We start by considering the general case. Using the same notation as in Section B.3, suppose that the first $d$ elements of $x$ and $\tilde{x}$ differ, i.e., $x_{i}\neq\tilde{x}_{i}$ for $i=1,\ldots,d$ and $x_{j}=\tilde{x}_{j}$ for $j=d+1,\ldots,n$. Without loss of generality, assume that $\sum_{i=1}^{d}v_{i}\geq\sum_{i=1}^{d}\tilde{v}_{i}$. Now consider the following cases:
@@
    Case (b). $I=\tilde{I}\in\{d+1,\ldots,n\}$. This case can increase the coupling distance only when one chain accepts the update while the other rejects it. The probability of this occurring is given by $|\alpha_{I}-\tilde{\alpha}_{\tilde{I}}|$.
@@
Let =$\mathcal{I}=\{d+1,\ldots,n\}$ be the coupled index, then we have:
$\quad P_{\mathrm{SOMA-coupled}}\left(d(x^{\prime},\tilde{x}^{\prime})=d(x,\tilde{x})+1\mid x,\tilde{x}\right)$

### v2

@@ A.3 SOMA as a multi-proposal sampler
We present another proof of Theorem 9.
Proof.
SOMA simultaneously considers $n$ proposed states $x_{-1},y]$, $x_{-2},y],\ldots,[x_{-n},y]$, with (unnormalized) weights $w_{i}(y,x).$ In this view, SOMA is indeed a multi-proposal MCMC method. The Markov transition kernel of SOMA $P_{\mathrm{SOMA}}(\cdot\mid x)$ is a special case of the $P^{(K)}(\cdot\mid x)$ from Pozza and Zanella, (2025) with $K=n$. SOMA has a special way of constructing multiple proposals and a weight function to choose among the proposed states.
In SOMA, the ‘marginal proposal’ for each component $i$ is in fact the IMH proposal $f(\cdot\mid\theta)$. Let
@@
denote the IMH proposal kernel for the $i$-th component, defined on the full state space $\mathbb{X}^{n}.$ Then $\widetilde{Q}(\cdot\mid x)=\frac{1}{n}Q_{i}(\cdot\mid x)$ is the overall proposal kernel for Ran-IMwG: it proposes an index $i$ uniformly and proposes a new from $\mathbb{X}$. We denote the Markov transition kernel associated with proposal $\widetilde{Q}$ and acceptance probability (5) by $\widetilde{P}$. It coincides with the Markov transition kernel of Ran-IMwG, denoted $P_{\mathrm{RAN}}(\cdot\mid x).$ In other words, Ran-IMwG is the ‘single proposal’ kernel that we should compare SOMA with. Pozza and Zanella, (2025, Theorem 1) states that
@@
which is equivalent to (12) in Theorem 9. ∎
A.4 Asymptotic Variance of SOMA and Computation Cost
@@
Luckily, computing the weights in parallel can reduce the runtime cost of SOMA. Oftentimes, the target distribution $\pi$ has the structure to make this parallel/vectorized implementation feasible. SOMA can thus have a computational complexity comparable to Ran-IMwG and Sys-IMwG under moderate values of $n$, using a vectorized implementation. The costs would have the same order under a parallel implementation. These allow one to harness the efficiency gained from SOMA’s design. We give Example 4, which continues with the perturbed summary statistics privacy mechanism in Section 5.2.
@@
Privacy mechanisms that add data-independent noise to a query of the form $s(x)=\sum_{i=1}^{n}t(x_{i})$ have densities $\eta(s_{\mathrm{dp}}\mid s(x))=g(s_{\mathrm{dp}},\sum_{i=1}^{n}t_{i}(x_{i}))$ for tractable and known functions $g$ and $t$. Ju et al., (2022, Assumption 2) coined the term ‘record additivity’ to describe how $\eta(s_{\mathrm{dp}}\mid s(x))$ depends on an additive summary of $x$, and connects this property with implementation efficiency of data augmentation methods. With the widely used Laplace mechanism $s_{\mathrm{dp}}\sim\mathrm{Laplace}(s(x),{\Delta}/{\epsilon})$ and the weights $w_{i}=\eta(s_{\mathrm{dp}}\mid s([x_{-i},y]))$, we have
@@
Examples 1, 3 and 2 are all instances of record addivite privacy mechanisms.
The swap-one-out structure in (23) enables parallel or vectorized implementation, which can reduce SOMA’s computational complexity to a large extent. We include empirical run time comparisons in Section B.4, showing the runtime cost of SOMA is less than twice that of the Ran/Sys-IMwG methods with $n=100$ components. In this case, instead of comparing the asymptotic variance of using $T$ iterations of SOMA with $nT$ iterations of Ran-IMwG, we should use $2T$ iterations of Ran-IMwG.
A.5 Proofs of Section 4.1
@@
Proof of Theorem 9.
Suppose $x$ and $x^{\prime}$ differ on component $i$ and, without loss of generality, assume $i=1$. Then, starting from the right-hand side,
$\quad n\cdot P_{\mathrm{RAN}}(x^{\prime}\mid x)$
$=n\cdot\frac{1}{n}f(x^{\prime}_{1}\mid\theta)\alpha_{1}^{\mathrm{IMwG}}(x^{\prime}_{1},x)=n\cdot\frac{1}{n}f(x^{\prime}_{1}\mid\theta)\min\left(\frac{w_{1}}{w_{0}},1\right)$
$=f(x^{\prime}_{1}\mid\theta)\cdot w_{1}\cdot\min\left(\frac{1}{w_{1}},\frac{1}{w_{0}}\right)$
$\geq f(x^{\prime}_{1}\mid\theta)\cdot w_{1}\cdot\min\left(\frac{1}{w_{1}+\sum_{i=2}^{n}w_{i}},\frac{1}{w_{0}+\sum_{i=2}^{n}w_{i}}\right)$
$=f(x^{\prime}_{1}\mid\theta)\cdot\frac{w_{1}}{w_{1}+\sum_{i=2}^{n}w_{i}}\min\left(1,\frac{w_{1}+\sum_{i=2}^{n}w_{i}}{w_{0}+\sum_{i=2}^{n}w_{i}}\right)$
$=P_{\mathrm{SOMA}}(x^{\prime}\mid x).$
@@
Based on Lemma 16, integrating both sides over $f(y\mid\theta)$ yields
$\int_{\mathbb{X}}f(y\mid\theta)\sum_{i=1}^{n}\left(\frac{w_{i}(y,x)}{W(y,x)}\right)\cdot\alpha_{i}^{\mathrm{SOMA}}(y,x)\mathrm{d}y\geq\int_{\mathbb{X}}f(y\mid\theta)\sum_{i=1}^{n}\left(\frac{1}{n}\right)\cdot\alpha_{i}^{\mathrm{IMwG}}(y,x)\mathrm{d}y,$
@@
Since $\varphi(u)=\mathbb{I}_{x}(u)$, the integral reduces to $1-A(x)]\cdot\int_{\mathbb{X}^{n}}\varphi(u)\mathrm{d}P^{m-1}(u\mid x)$. By induction, we have
@@
We first prove that $A^{\mathrm{SOMA}}(x)\geq{n}/({n+e^{\epsilon}-1})$. Let $T_{i}(y,x)=\sum_{j=1,j\neq i}^{n}w_{j}(y,x)$. Under the definition of $\epsilon$-DP, we have $w_{i}(y,x)/w_{0}(y,x)\geq 1/e^{\epsilon}$, and $T_{i}(y,x)/w_{0}(y,x)\geq(n-1)/e^{\epsilon}$ holds for all $x$ and $y$. Now the overall acceptance probability $A^{\mathrm{SOMA}}(x)$ for SOMA has the form
$A^{\mathrm{SOMA}}(x)$ $=\int_{\mathbb{X}}f(y\mid\theta)\sum_{i=1}^{n}\left[\mathbb{P}(I=i)\min\left\{1,\frac{W(y,x)}{W(y,x)+w_{0}(y,x)-w_{i}(y,x)}\right\}\right]\mathrm{d}y$
$=\int_{\mathbb{X}}f(y\mid\theta)\sum_{i=1}^{n}\left[\mathbb{P}(I=i)\min\left\{1,\frac{T_{i}(y,x)+w_{i}(y,x)}{T_{i}(y,x)+w_{0}(y,x)}\right\}\right]\mathrm{d}y$
$=\int_{\mathbb{X}}f(y\mid\theta)\sum_{i=1}^{n}\left[\mathbb{P}(I=i)\min\left\{1,1-\frac{1-w_{i}(y,x)/w_{0}(y,x)}{1+T_{i}(y,x)/w_{0}(y,x)}\right\}\right]\mathrm{d}y$
$\geq\int_{\mathbb{X}}f(y\mid\theta)\sum_{i=1}^{n}\left[\mathbb{P}(I=i)\min\left\{1,1-\frac{1-1/e^{\epsilon}}{1+(n-1)/e^{\epsilon}}\right\}\right]\mathrm{d}y$
$=1-\frac{1-1/e^{\epsilon}}{1+(n-1)/e^{\epsilon}}$
$=\frac{n}{n+e^{\epsilon}-1}.$ (34)
@@
$A^{\mathrm{RAN}}(x)$ $=\int_{\mathbb{X}}f(y\mid\theta)\sum_{i=1}^{n}\left[\mathbb{P}(I=i)\min\left\{1,\frac{w_{i}(y,x)}{w_{0}(y,x)}\right\}\right]\mathrm{d}y$
$\geq\int_{\mathbb{X}}f(y\mid\theta)\sum_{i=1}^{n}\left[\mathbb{P}(I=i)\min\left\{1,\frac{1}{e^{\epsilon}}\right\}\right]\mathrm{d}y$
$=\frac{1}{e^{\epsilon}},$ (35)
@@
$A_{i}^{\mathrm{SYS}}(x)$ $=\int_{\mathbb{X}}f(y\mid\theta)\min\left\{1,\frac{w_{i}(y,x)}{w_{0}(y,x)}\right\}\mathrm{d}y$
$\geq\int_{\mathbb{X}}f(y\mid\theta)\min\left\{1,\frac{1}{e^{\epsilon}}\right\}\mathrm{d}y$
$=\frac{1}{e^{\epsilon}}.$ (36)
@@
A.6 The Coupling Kernel of SOMA in Section 4.2
We now use coupling-based techniques to analyze the convergence rate of the algorithms. Our proof leverages the coupling lemma. It entails constructing two chains on a common probability space such that they eventually meet and remain coupled.
Theorem 17 (Coupling inequality (Rosenthal,, 1995)).
$\|P^{t}(\cdot\mid x)-\pi(\cdot)\|_{\mathrm{TV}}\leq\mathbb{P}(\tau>t).$ (37)
where the meeting time is defined by $\tau=\min\{t:\ \Phi^{(t)}=\tilde{\Phi}^{(t)}\}$.
To utilize (37), we must construct such a faithful coupling. We create a coupling of SOMA in the two stages: (1) use a common proposed state $y$ and select a common index $I$ using a maximal coupling of Multinomial random variables (Thorisson,, 2000; Jacob,, 2020); (2) accept and reject the proposed states with a common random number.
Assume we have two chains $\Phi$ and $\tilde{\Phi}$ with corresponding states $x^{(t)}$ and $\tilde{x}^{(t)}$ at time $t$. Here is a coupling kernel.
1. Propose the same single offer $y\sim f(y\mid\theta)$ for both chains $\Phi$ and $\tilde{\Phi}$.
@@ 3. Compute the component-wise minimum of selection probabilities and residuals:
[…]
@@ 4. Calculate the acceptance probability
We now provide the following property of the coupling kernel.
Lemma 18.
For any current state (x,x~)($\tilde{x}$), let (x,x~)($^{\prime},\tilde{x}^{\prime}$) be samples from the coupled SOMA kernel $P_{\mathrm{SOMA-coupled}}$. Under $\epsilon$-DP, we must have
[…]
$\geq\frac{n}{n+M-1}\cdot\frac{d(x,\tilde{x})}{d(x,\tilde{x})+e^{\epsilon}(n-d(x,\tilde{x}))}$. (40)
$P_{\mathrm{Ran-coupled}}\left(d(x^{\prime},\tilde{x}^{\prime})=d(x,\tilde{x})-1\mid x,\tilde{x}\right)\geq\frac{1}{e^{\epsilon}}\cdot\frac{d(x,\tilde{x})}{n}.$ (41)
@@
Let the set of coupled indices be $\mathcal{I}=\{i:x_{i}=\tilde{x}_{i}\}$. In the coupling construction, if $I$ and $\tilde{I}$ are both on uncoupled positions and both chains accept the proposed component $y$, then the distance between the chains will reduce by 1. This argument leads to
[…]
$=\int_{\mathbb{X}}\sum_{i\in\mathcal{I}^{c}}\sum_{j\in\mathcal{I}^{c}}\mathbb{P}(I=i,\tilde{I}=j\mid x,\tilde{x},y)\cdot\min(\alpha_{i},\tilde{\alpha}_{j};x,\tilde{x},y)f(y\mid\theta)\mathrm{d}y$
$\geq\frac{n}{n+e^{\epsilon}-1}\cdot\mathbb{P}(I\in\mathcal{I}^{c},\tilde{I}\in\mathcal{I}^{c}\mid x,\tilde{x},y)$
$=\frac{n}{n+e^{\epsilon}-1}\cdot\min\left(\frac{\sum_{i\in\mathcal{I}^{c}}w_{i}}{\sum_{i=1}^{n}w_{i}},\frac{\sum_{i\in\mathcal{I}^{c}}\tilde{w_{i}}}{\sum_{i=1}^{n}\tilde{w}_{i}}\right)$
$\geq\frac{n}{n+e^{\epsilon}-1}\cdot\frac{d(x,\tilde{x})}{d(x,\tilde{x})+e^{\epsilon}(n-d(x,\tilde{x}))}$.
The first inequality uses Theorem 12. The second lower bound is achieved when ${w_{j}}/{w_{i}}=e^{\epsilon}$ for all $j\in\mathcal{I}$ and all $i\in\mathcal{I}^{c}$. We also note that the two index sets have cardinality d(x,x~)d($\tilde{x}$) and nd(x,x~)n-d($\tilde{x}$) respectively.
@@
$P_{\mathrm{Ran-coupled}}\left(d(x^{\prime},\tilde{x}^{\prime})=d(x,\tilde{x})-1\mid x,\tilde{x}\right)\geq\inf_{i\in\mathcal{I}^{c}}\left[\alpha_{i}^{\mathrm{IMwG}}(x,y)\right]\cdot\mathbb{P}(I=\tilde{I}\in\mathcal{I}^{c})=\frac{1}{e^{\epsilon}}\cdot\frac{d(x,\tilde{x})}{n}.$
@@
For any current state (x,x~)($\tilde{x}$), let (x,x~)($^{\prime},\tilde{x}^{\prime}$) be samples from the coupled SOMA kernel $P_{\mathrm{SOMA-coupled}}$. Under $\epsilon$-DP, we must have
[…]
$\leq\frac{(n-d)e^{\epsilon}}{(n-d)e^{\epsilon}+d}-\frac{n}{n+e^{\epsilon}-1}\cdot\frac{(n-d)}{(n-d)+de^{\epsilon}},$ (42)
where we use d=d(x,x~)d=d($\tilde{x}$) for simplicity.
@@
$P_{\mathrm{SOMA-coupled}}\left(d(x^{\prime},\tilde{x}^{\prime})=d(x,\tilde{x})+1\mid x,\tilde{x}\right)\leq\frac{e^{\epsilon}(e^{\epsilon}-1)}{(e^{\epsilon}+1)^{2}}.$ (43)
@@
We start by considering the general case. Using the same notation as in Section A.6, suppose that the first $d$ elements of $x$ and $\tilde{x}$ differ, i.e., $x_{i}\neq\tilde{x}_{i}$ for $i=1,\ldots,d$ and $x_{j}=\tilde{x}_{j}$ for $j=d+1,\ldots,n$. Without loss of generality, assume that $\sum_{i=1}^{d}v_{i}\geq\sum_{i=1}^{d}\tilde{v}_{i}$. Now consider the following cases:
@@
[…]
@@
Let $\mathcal{I}=\{d+1,\ldots,n\}$ be the coupled index, then we have:
@@
$=\int_{\mathbb{X}}\sum_{i\in\mathcal{I}}\mathbb{P}(I=\tilde{I}=i\mid x,\tilde{x},y)\cdot(|\alpha_{i}-\tilde{\alpha}_{j}|;x,\tilde{x},y)f(y\mid\theta)\mathrm{d}y\quad(\text{Case (b)})$
$\quad+\int_{\mathbb{X}}\sum_{i\in\mathcal{I}^{c}}\sum_{j\in\mathcal{I}}\mathbb{P}(I=i,\tilde{I}=j\mid x,\tilde{x},y)\cdot(\max\{\tilde{\alpha}_{j}-\alpha_{i},0\};x,\tilde{x},y)f(y\mid\theta)\mathrm{d}y\quad(\text{Case (d)})$
$\quad+\int_{\mathbb{X}}\sum_{i\in\mathcal{I}}\sum_{j\in\mathcal{I}}\mathbb{P}(I=i,\tilde{I}=j,i\neq j\mid x,\tilde{x},y)\cdot(\max\{\alpha_{i},\tilde{\alpha}_{j}\};x,\tilde{x},y)f(y\mid\theta)\mathrm{d}y.\quad(\text{Case (f)})$ (44)
@@
$\leq\int_{\mathbb{X}}\sum_{i=d+1}^{n}u_{i}\cdot(|\alpha_{i}-\tilde{\alpha}_{i}|;x,\tilde{x},y)f(y\mid\theta)\mathrm{d}y+\int_{\mathbb{X}}\sum_{i=d+1}^{n}\tilde{v}_{i}\cdot 1\cdot f(y\mid\theta)\mathrm{d}y$
$\leq\int_{\mathbb{X}}\left[\left(1-\frac{n}{n+e^{\epsilon}-1}\right)\sum_{i=d+1}^{n}u_{i}+\sum_{i=d+1}^{n}\tilde{v}_{i}\right]f(y\mid\theta)\mathrm{d}y$
$=\int_{\mathbb{X}}\sum_{i=d+1}^{n}(\tilde{v}_{i}+u_{i})f(y\mid\theta)\mathrm{d}y-\frac{n}{n+e^{\epsilon}-1}\int_{\mathbb{X}}\sum_{i=d+1}^{n}u_{i}f(y\mid\theta)\mathrm{d}y$
$=\int_{\mathbb{X}}\frac{\sum_{i=d+1}^{n}\tilde{w}_{i}}{\tilde{W}}f(y\mid\theta)\mathrm{d}y-\frac{n}{n+e^{\epsilon}-1}\int_{\mathbb{X}}\sum_{i=d+1}^{n}u_{i}f(y\mid\theta)\mathrm{d}y.$ (45)
Now, under the $\epsilon$-DP, $\frac{\sum_{i=d+1}^{n}w_{i}}{W}$ has the maximum value $\frac{(n-d)e^{\epsilon}}{(n-d)e^{\epsilon}+d}$ when $w_{1}=w_{2}=\ldots=w_{d}$ and $e^{\epsilon}w_{1}=w_{d+1}=w_{d+2}=\ldots=w_{n}$. As for $\frac{\sum_{i=d+1}^{n}w_{i}}{W}$, it has the minimum value $\frac{(n-d)}{(n-d)+de^{\epsilon}}$ when $\tilde{w}_{1}=\tilde{w}_{2}=\ldots=\tilde{w}_{d}$ and $\tilde{w}_{1}/e^{\epsilon}=\tilde{w}_{d+1}=\tilde{w}_{d+2}=\ldots=\tilde{w}_{n}$. In this case, $\sum_{i=d+1}^{n}u_{i}=\sum_{i=d+1}^{n}\min\left\{\frac{w_{i}}{W},\frac{\tilde{w_{i}}}{\tilde{W}}\right\}$ has the minimun value $\frac{(n-d)}{(n-d)+de^{\epsilon}}$. It then gives
$P_{\mathrm{SOMA-coupled}}\left(d(x^{\prime},\tilde{x}^{\prime})d(x,\tilde{x})+1\mid x,\tilde{x}\right)\leq\frac{(n-d)e^{\epsilon}}{(n-d)e^{\epsilon}+d}-\frac{n}{n+e^{\epsilon}-1}\cdot\frac{(n-d)}{(n-d)+de^{\epsilon}}.$
For the case $n=2$ and $d(x,\tilde{x})=1$, Case (f) cannot occur and we have